OpenStudy (math2400):

radius of convergence and the interval of convergence for the given power series

OpenStudy (math2400):

\[\sum_{n=1}^{\infty} \frac{ n ^{2}x ^{2n} }{ n! }\]

OpenStudy (math2400):

my radius is 1. And it think the interval is (-1,1) Can anyone verify if that's right?

OpenStudy (zarkon):

what test did you use?

OpenStudy (math2400):

ratio

OpenStudy (zarkon):

do it again

OpenStudy (math2400):

ok

OpenStudy (math2400):

wait can the radius be infinity??

OpenStudy (zarkon):

yes

OpenStudy (math2400):

ok..and the interval is (-infin, infin) ?

OpenStudy (zarkon):

yes

OpenStudy (math2400):

lol kk thanks. that's what i got first but i wasn't sure if that was possible haha

OpenStudy (xapproachesinfinity):

i got oo as well :)

OpenStudy (math2400):

perfect thanks!! :)

OpenStudy (xapproachesinfinity):

limit (x^2 . (n+1/n)^2/(n+1))=0

OpenStudy (math2400):

how about this one if u don't mind: \[\sum_{n=1}^{\infty} \frac{ n^{e} }{ e ^{n} }\] i'm pretty sure the radius is 1 right?

OpenStudy (xapproachesinfinity):

that says for any value of x rho is always zero

OpenStudy (math2400):

and then interval is (2,4) or maybe (2,4] not sure, i suck at these

OpenStudy (xapproachesinfinity):

i would do root test with that one

OpenStudy (xapproachesinfinity):

did you try ratio?

OpenStudy (zarkon):

there is no x in your last problem

OpenStudy (math2400):

yes(: I got the radius to be 1 and then i know the interval is (2,4) but I'm not sure

OpenStudy (xapproachesinfinity):

eh hold on you are asking for convergence because there is not x

OpenStudy (xapproachesinfinity):

you need x

OpenStudy (math2400):

ohh wait sorry i posted the wrong one. That one is a converging series haha. I meant to post this one: \[\sum_{n=1}^{\infty} \frac{ (x-3)^n }{ \sqrt{n} }\]

OpenStudy (xapproachesinfinity):

oh ok i would go with root test with that too limit n^1/n=1 rho=|x-3|<1 so raduis is 1 -1<x-3<1 ===> 2<x<4

OpenStudy (xapproachesinfinity):

i guess you got this :)

OpenStudy (math2400):

ok so radius is 1 and then did u check the end points?

OpenStudy (xapproachesinfinity):

No i didn't x=4 you have series 1/n^1/2 diverges x=2 (-1)^n/rootn i believe this converge by alternating series test

OpenStudy (math2400):

so would it be (2,4] or just (2,4) that's what I'm confused on

OpenStudy (xapproachesinfinity):

no converge on 2 not on 4 so [2,4) otherwise 2<=x<4

OpenStudy (math2400):

ohh haha kk got it :) Thank you so much!

OpenStudy (xapproachesinfinity):

no problem :)

OpenStudy (xapproachesinfinity):

actually you seem to me good at this, just few more practice and you master it

OpenStudy (math2400):

haha thank you, means a lot @xapproachesinfinity